OpenStudy (anonymous):

Rational root x^3−2x^2+10x+136=0

OpenStudy (nikvist):

\[x^3-2x^2+10x+136=0\] \[x^3+4x^2-6x^2-24x+34x+136=0\] \[(x+4)(x^2-6x+34)=0\] \[x_1=-4\quad,\quad x_{2,3}=\frac{6\pm\sqrt{36-136}}{2}=3\pm 5i\]